# Reassessing How to Compare and Improve the Calibration of Machine Learning Models

### v1

@@ Abstract
A machine learning model is calibrated if its predicted probability for an outcome matches the observed frequency for that outcome conditional on the model prediction. This property has become increasingly important as the impact of machine learning models has continued to spread to various domains. As a result, there are now a dizzying number of recent papers on measuring and improving the calibration of (specifically deep learning) models. In this work, we reassess the reporting of calibration metrics in the recent literature. We show that there exist trivial recalibration approaches that can appear seemingly state-of-the-art unless calibration and prediction metrics (i.e. test accuracy) are accompanied by additional generalization metrics such as negative log-likelihood. We then derive a calibration-based decomposition of Bregman divergences that can be used to both motivate a choice of calibration metric based on a generalization metric, and to detect trivial calibration. Finally, we apply these ideas to develop a new extension to reliability diagrams that can be used to jointly visualize calibration as well as the estimated generalization error of a model.
@@ 1.1 Summary of Main Contributions and Takeaways
    We identify multiple issues in how calibration performance is reported in the context of multi-class classification, with the core issue being that many (even very recent) papers report only some notion(s) of calibration error along with test accuracy. With respect to these metrics, the trivial recalibration strategy of always predicting with the mean confidence obtained from a held-out calibration set outperforms the most popular post-training recalibration methods.
@@
    Based on this observation, we revisit the use of proper scoring rules for measuring calibration. We focus on the particular case of scoring rules that are Bregman divergences, and derive a decomposition of the divergence into a calibration error and a leftover “sharpness” term. This latter term can be thought of as corresponding to how much information is preserved by a predictor; i.e. a constant predictor preserves no information from the original data while an injective predictor preserves everything. We show how our decomposition can be used to detect cases of “trivial calibration” and avoid potentially misleading results, and along the way prove several new results connecting Bregman divergences and notions of calibration in the literature.
@@
    Finally, we propose an extension to reliability diagrams based on the aforementioned decomposition that can be used to jointly visualize calibration and sharpness, and show that our visualization can be used to quickly discern important trade-offs between recalibration methods. We also prove the consistency of our estimation approach, thereby extending existing estimation results. Our visualization approach is available as an open-source Python package, and can be installed simply via pip install sharpcal or from source at https://github.com/2014mchidamb/reassessing-calibration.
The main takeaway of our results is that the choice of calibration metric should be motivated by the choice of generalization metric or loss function and that both should always be reported together in order to ensure that improvements in model calibration are not at the cost of model sharpness.
@@ 1.2 Related Work
Calibration and proper scoring rules. Divergences obtained from strictly proper scores such as negative log-likelihood (NLL) and mean-squared error (MSE) often serve as metrics for quantifying generalization. A predictor that is optimal with respect to a strictly proper scoring rule has recovered the ground-truth conditional distribution, and is therefore necessarily calibrated. However, requiring optimality with respect to a proper scoring rule is a stronger condition than requiring calibration, and several works have thus explored how to decompose proper scoring rules into a calibration term and a sharpness/resolution term (Murphy, 1973; Gneiting et al., 2007; Bröcker, 2008; Kull & Flach, 2015; Pohle, 2020). These works are either at the specificity of a particular scoring rule (Murphy, 1973) or at the generality of strictly proper scoring rules satisfying certain consistency conditions (Pohle, 2020). In this work we focus on the in-between regime of Bregman divergences, for which we can use the additional structure to prove novel connections to the calibration definitions in the machine learning literature.
@@ 3 Pitfalls in Calibration Reporting
The ramifications of the MRR strategy seem to be under-appreciated in the literature - we have minimized calibration error without affecting the prediction of $g$ by using a constant confidence function $h$. By separating the confidence function from the classification function, we have made it so that the test accuracy remains unchanged, so that this kind of trivial confidence behavior (trivial in a sense that is made rigorous in Section 4) is undetectable by just reporting calibration error and accuracy.
@@
The above listing of papers is of course non-exhaustive; we have focused on a handful of influential papers (all have at least 100 citations) to merely point out that reporting (or at least focusing on) only ECE/similar and test accuracy is common practice. That being said, we are not at all calling into question the advances made by the above papers - many of their findings have been repeatedly corroborated by follow-up work. On the other hand, for new work on calibration we strongly advocate for always reporting errors derived from proper scoring rules (i.e. NLL or MSE) alongside pure calibration metrics such as ECE (we put forth a methodology for doing so in Sections 4 and 5).
Experiments. To drive home the point, we revisit the experiments of Guo et al. (2017), but focus on one of the more modern architectural choices of Minderer et al. (2021). Namely, we evaluate the most downloaded pretrained vision transformer (ViT) model [^1] The model card is available at: https://huggingface.co/timm/vit_base_patch16_224.augreg2_in21k_ft_in1k. (Dosovitskiy et al., 2021; Steiner et al., 2021) available through the timm library (Wightman, 2019) with respect to binned ECE, binned ACE (Nixon et al., 2019), SmoothECE (Błasiok & Nakkiran, 2023), negative log-likelihood, and MSE on ImageNet-1K-Val (Russakovsky et al., 2015). We split the data into 10,000 calibration samples (20% split) and 40,000 test samples, and compare the unmodified test performance to temperature scaling (TS), histogram binning (HB), isotonic regression (IR), and our proposed MRR. We follow the same TS implementation as Guo et al. (2017) and also use 15 bins for the binning estimators (ECE, ACE, HB) to be comparable to the results in their work.
@@
4 Calibration-Sharpness Decomposition for Bregman Divergences
As can be seen in Table 1, the MRR strategy dominates in terms of all of the pure calibration metrics, but leads to large jumps in NLL and MSE. In the specific case of Table 1, either NLL or MSE alone would have been sufficient to detect the trivial behavior of MRR. We now address the more general question: which additional generalization metrics should we report to prevent misleading calibration results? Our answer relies on the fact that such metrics are usually Bregman divergences, and we prove a decomposition of Bregman divergences that relates calibration and predictive power.
4.1 General Theory
Let us use $\phi:\Delta^{k-1}\to\overline{\mathbb{R}}$ to denote a continuously differentiable, strictly convex function. We recall that the Bregman divergence $d_{\phi}$ associated with $\phi$ is defined to be:
@@ 4.1 General Theory
In many machine learning tasks the goal is to minimize the expectation of a conditional Bregman divergence; for example, in multi-class classification we take $\phi(x)=\sum_{i}x_{i}\log x_{i}$ to obtain the KL divergence and then attempt to minimize $\mathbb{E}[d_{\phi}(\pi_{Y\mid X},g(X))]$ from samples. We now show that every $d_{\phi}$ can be decomposed into a calibration error term and what we refer to as a sharpness gap term.

### v2

@@ Abstract
A machine learning model is calibrated if its predicted probability for an outcome matches the observed frequency for that outcome conditional on the model prediction. This property has become increasingly important as the impact of machine learning models has continued to spread to various domains. As a result, there are now a dizzying number of recent papers on measuring and improving the calibration of (specifically deep learning) models. In this work, we reassess the reporting of calibration metrics in the recent literature. We show that there exist trivial recalibration approaches that can appear seemingly state-of-the-art unless calibration and prediction metrics (i.e. test accuracy) are accompanied by additional generalization metrics such as negative log-likelihood. We then use a calibration-based decomposition of Bregman divergences to develop a new extension to reliability diagrams that jointly visualizes calibration and generalization error, and show how our visualization can be used to detect trade-offs between calibration and generalization. Along the way, we prove novel results regarding the relationship between full calibration error and confidence calibration error for Bregman divergences. We also establish the consistency of the kernel regression estimator for calibration error used in our visualization approach, which generalizes existing consistency results in the literature.
@@ 1.1 Summary of Main Contributions and Takeaways
    We identify multiple issues (identified in Sections 3, 4.1, and 4.2) in how calibration performance is reported in the context of multi-class classification, with the core issue being that many (even very recent) papers report only some notion(s) of calibration error along with test accuracy. With respect to these metrics, the trivial recalibration strategy of always predicting with the mean confidence obtained from a calibration set outperforms the most popular post-training recalibration methods.
@@
    Based on this observation, we revisit the use of proper scoring rules for measuring calibration. We apply decomposition results for these scoring rules to show how we can motivate a calibration metric based on our choice of generalization metric, and in special cases how we can do the converse as well. Furthermore, we focus on proper scoring rules that are Bregman divergences and leverage the structure of these divergences to also prove new results regarding the relationship between full calibration and confidence calibration, with our main result being Lemma 4.6 (confidence calibration error is a lower bound for full calibration error).
@@
    Finally, we propose an extension to reliability diagrams for jointly visualizing calibration and generalization. We also prove the consistency of the estimation approach we use in our new visualization in Theorem 4.9, thereby extending existing estimation results in the literature. In Section 5, we show how our calibration-sharpness diagrams can provide a more granular comparison of recalibration methods when compared to traditional reliability diagrams (Figure 2). Code for our visualizations is available as the Python package sharpcal.
The main takeaway of our results is that the choice of calibration metric should be motivated by the choice of generalization metric or loss function and that both should be reported together in order to ensure that improvements in model calibration are not at the cost of model sharpness.
@@ 1.2 Related Work
Calibration and proper scoring rules. Divergences obtained from strictly proper scores such as negative log-likelihood (NLL) and mean-squared error (MSE) often serve as metrics for quantifying generalization. A predictor that is optimal with respect to a strictly proper scoring rule has recovered the ground-truth conditional distribution, and is therefore necessarily calibrated. However, requiring optimality with respect to a proper scoring rule is a stronger condition than requiring calibration, and several works have thus explored how to decompose proper scoring rules into a calibration term and a sharpness/resolution term (Murphy, 1973; Gneiting et al., 2007; Bröcker, 2008; Kull & Flach, 2015; Pohle, 2020). The most related such work to ours is that of Perez-Lebel et al. (2023), which also puts forth the perspective that reporting calibration and accuracy is insufficient; however, they focus on a decomposition involving a grouping loss term whereas we focus on a sharpness term which leads to different visualizations and different takeaways. We provide a more detailed comparison to this work in Appendix E.
@@ 3 Pitfalls in Calibration Reporting
The ramifications of the MRR strategy seem to be under-appreciated in the literature - we have minimized calibration error without affecting the prediction of $g$ by using a constant confidence function $h$. By separating the confidence function from the classification function, we have made it so that the test accuracy remains unchanged, so that this kind of trivial confidence behavior is undetectable by just reporting calibration error and accuracy. Even though MRR may seem extreme, one can imagine modifications where we effectively perform the procedure of replacing the predicted confidence with the calibration accuracy while doing something slightly more clever with the other predicted probabilities.
@@
The above listing of papers is of course non-exhaustive; we have focused on a handful of influential papers (all have at least 100 citations) to merely point out that reporting (or at least focusing on) only ECE/similar and test accuracy is common practice. That being said, we are not at all calling into question the advances made by the above papers - many of their findings have been repeatedly corroborated by follow-up work. On the other hand, for new work on calibration we strongly advocate for always reporting errors derived from proper scoring rules (i.e. NLL or MSE) alongside pure calibration metrics such as ECE (we put forth a methodology for doing so in Sections 4.1 and 4.2).
Experiments. To drive home the point, we revisit the experiments of Guo et al. (2017), but focus on one of the more modern architectural choices of Minderer et al. (2021). Namely, we evaluate the most downloaded pretrained vision transformer (ViT) model [^1] The model card is available at: https://huggingface.co/timm/vit_base_patch16_224.augreg2_in21k_ft_in1k. (Dosovitskiy et al., 2021; Steiner et al., 2021) available through the timm library (Wightman, 2019) with respect to binned ECE, binned ACE (Nixon et al., 2019), SmoothECE (Błasiok & Nakkiran, 2023), negative log-likelihood, and MSE on ImageNet-1K-Val (Russakovsky et al., 2015). We split the data into 10,000 calibration samples (20% split) and 40,000 test samples, and compare the unmodified test performance to temperature scaling (TS), histogram binning (HB), isotonic regression (IR), and our proposed MRR. We follow the same TS implementation as Guo et al. (2017) and also use 15 bins for the binning estimators (ECE, ACE, HB) to be comparable to the results in their work. The results shown in Table 1 are not sensitive to the choice of model; we show similar results for several other popular timm models in Appendix F.
@@
4 Main Theoretical Results
As can be seen in Table 1, the MRR strategy dominates in terms of all of the pure calibration metrics, but leads to large jumps in NLL and MSE. Clearly either NLL or MSE alone would have been sufficient to detect the trivial behavior of MRR. We now address the more general question: which generalization metrics should we report to prevent misleading calibration results? Our answer relies on the fact that such metrics are usually Bregman divergences. Our focus on Bregman divergences is due to their convenient theoretical properties; in particular, they are defined in terms of convex functions and we can thus bring to bear the machinery of convex analysis.
We first review the definitions. Let us use $\phi:\Delta^{k-1}\to\overline{\mathbb{R}}$ to denote a continuously differentiable ($C^{1}$), strictly convex function. We recall that the Bregman divergence $d_{\phi}$ associated with $\phi$ is defined to be:
@@ 4 Main Theoretical Results
In many machine learning tasks the goal is to minimize the expectation of a conditional Bregman divergence; for example, in multi-class classification we take $\phi(x)=\sum_{i}x_{i}\log x_{i}$ to obtain the KL divergence and then attempt to minimize $\mathbb{E}[d_{\phi}(\pi_{Y\mid X},g(X))]$ from samples. Bregman divergences are closely related with strictly proper scoring rules, and we can leverage the fact that such rules can be decomposed into a calibration error and a “sharpness” term to obtain a similar decomposition for Bregman divergences.
@@
More general versions of Lemma 4.1 have appeared in Bröcker (2009); Kull & Flach (2015); Pohle (2020); Gruber & Buettner (2024), but we present a version for Bregman divergences as we will leverage the structure of (4.1) to prove our new results in the subsequent sections. We provide a self-contained proof of Lemma 4.1 in Appendix A.
Some intuition for the sharpness gap term is provided in the next two propositions, the first of which appeared in Pohle (2020) in a different form. The second follows easily from the first and Lemma 4.1, but has not appeared elsewhere to our knowledge. Intuitively, the propositions show that a smaller sharpness gap corresponds to a predictor that is more “fine-grained”, and in that sense the constant calibrated predictor that we introduced in the previous subsection is the worst possible fully calibrated predictor. The proofs of both propositions can once again be found in Appendix A.
Proposition 4.2.
For $g_{1},g_{2}:\mathbb{R}^{d}\to\Delta^{k-1}$ and $\pi$ the joint distribution of data and label $(X,Y)$ on $\mathbb{R}^{d}\times[k]$, if $\sigma(g_{1}(X))\subseteq\sigma(g_{2}(X))$, then:
Proposition 4.3.
The main takeaway from Lemma 4.1 is that when we choose to optimize a Bregman divergence, we implicitly also have a notion of calibration error that we are minimizing via (4.2). This provides a practical workflow to answer our motivating question of which metrics to report together: we first fix a Bregman divergence for our notion of generalization, and then report the same divergence but applied to the expected label conditioned on the predicted probabilities. We can also ask about the other direction – if we choose a calibration error, is there a natural Bregman divergence associated with it?
The answer is: not always. In fact, in considering this question, we can motivate a choice of $L^{q}$ ECE. The next result shows that in the framework of Bregman divergences the “right” choice of ECE is the $L^{2}$ ECE, since it is the conditional expectation of the Brier score.
Proposition 4.4.
Let $\pi$ correspond to the joint distribution of data and label $(X,Y)$ on $\mathbb{R}^{d}\times[k]$, and define the $L^{q}$ (for integers $q\geq 1$) ECE of a function $g:\mathbb{R}^{d}\to\Delta^{k-1}$ to be $\mathbb{E}\left[\norm{\mathbb{E}[Y\mid g(X)]-g(X)}_{q}^{q}\right]$. Then only the $L^{2}$ ECE is induced by a Bregman divergence in the sense of Lemma 4.1.
The proof is a simple consequence of characterizations provided in Boissonnat et al. (2010) (which we recall in Appendix A).
4.1 Confidence Calibration With Bregman Divergences
So far, we have leveraged existing ideas in the literature on proper scoring rules to motivate a joint choice of generalization and calibration error. However, as we alluded to earlier, when considering high-dimensional classification problems it is not feasible to efficiently estimate the conditional expectation required for calibration error. The main result of this section is to show that we can replace the full calibration error in (4.2) with confidence calibration error and still obtain a valid lower bound, which provides justification for the reporting of confidence calibration at least in the context of Bregman divergences.
In order to discuss confidence calibration, we need a 1-D counterpart $\varphi:[0,1]\to\overline{\mathbb{R}}$ to $\phi$. For our cases of interest, the choice of $\varphi$ is obvious; for example, when $\phi(x)=\norm{x}^{2}$ (i.e. $d_{\phi}$ is Brier score), $\varphi(x)=x^{2}$. Recalling that in standard confidence calibration we have $c(g(X))=\operatorname*{argmax}_{i}g^{i}(X)$ and $h(g(X))=\max_{i}g^{i}(X)$, our goal is then to investigate the relationship between $\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X))]$ and $\mathbb{E}[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))]$.
Prior to doing so, let us pause to stress the fact that the confidence calibration framework should be applied only to the calibration error and not to the entire loss. In other words, we should report the full Bregman divergence $\mathbb{E}[d_{\phi}(\pi_{Y\mid X},g(X))]$ in addition to the calibration error, not the 1-D Bregman divergence $\mathbb{E}[d_{\varphi}(\mathbbm{1}_{Y=c(X)},h(X))]$. Reporting of the latter is another pitfall and has appeared in the recent literature on surveying confidence elicitation strategies for LLMs (Tian et al., 2023). We show in the next proposition that the 1-D loss can have very different behavior when compared to the full loss for the specific case of MSE/Brier score.
Proposition 4.5.
The proof follows from directly evaluating (4.5), and can be found in Appendix B. Proposition 4.5 shows that the loss $d_{\varphi}$ incurred on the confidence calibration problem can be arbitrarily small for a predictor that is always wrong despite such a predictor having true loss on the multi-class classification problem bounded below by a constant. We now return to our original goal, which was showing that $\mathbb{E}[d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X))]$ is a lower bound for $\mathbb{E}[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))]$ under the condition that $\phi$ dominates $\varphi$ applied to any individual coordinate pair (for example, this is easily seen to be the case for Brier score).
Lemma 4.6.
For any $C^{1}$, convex functions $\phi:\Delta^{k-1}\to\overline{\mathbb{R}}$ and $\varphi:[0,1]\to\overline{\mathbb{R}}$ such that their Bregman divergences satisfy $d_{\phi}(x,y)\geq d_{\varphi}(x_{i},y_{i})$ for all $i\in[k]$, it follows (with probability 1) that:
$\mathbb{E}[d_{\phi}(\mathbb{E}[\pi_{Y\mid X}\mid g(X)],g(X))\mid h(X)]\geq d_{\varphi}(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)],h(X)).$ (4.6)
Corollary 4.7.
The inequality (4.6) holds for the choices $\phi(x)=\sum_{i}x_{i}\log x_{i}$ (KL divergence) and $\varphi(x)=x\log x+(1-x)\log(1-x)$ (logistic loss).
The proof of Lemma 4.6 relies on a careful analysis of the conditional expectation terms, and can also be found in Appendix B. Corollary 4.7 follows from the data processing inequality.
4.2 Calibration-Sharpness Diagrams
To summarize our observations thus far, we have recommended that calibration should be compared using the calibration metric associated with the Bregman divergence for which the models were optimized, and that both should be reported together. It remains to address how to similarly modify the visualization of (confidence) calibration, which is typically done via reliability diagrams that plot $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]$ for different values of $h(X)$. Our goals are three-fold: we want a visualization that is efficient to compute, makes it obvious whether one model is better calibrated and/or has better overall divergence than another model, and allows one to pinpoint the regions of the predicted probability space where model performance is weak. While there are a few recent works that try to make progress on these goals, we are not aware of a solution presented as a single diagram satisfying our desiderata (Dimitriadis et al., 2023; Gneiting & Resin, 2023).
The approach we take is to augment the plot of $\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)]$ with a surrounding band whose length $\rho(h(X))$ corresponds to the sharpness gap from Lemma 4.1 at $h(X)$, which we refer to as calibration-sharpness diagrams. For convenience, we will use $\rho(X)$ in place of $\rho(h(X))$ when $h$ is understood. The function $\rho:[0,1]\to\overline{\mathbb{R}}$ is the pointwise sharpness gap, and is formally defined as:
$\rho(p)=\mathbb{E}\left[d_{\phi}(\pi_{Y\mid X},g(X))\mid h(X)=p\right]-d_{\varphi}\left(\mathbb{E}[\mathbbm{1}_{Y=c(X)}\mid h(X)=p],p\right),$ (4.7)
where as in the previous subsection $d_{\phi}$ is the full Bregman divergence and $d_{\varphi}$ is its 1-D counterpart. Note that in (4.7) we are conditioning $d_{\phi}(\pi_{Y\mid X},g(X))$ on $h(X)$, since we want a value for the divergence at different confidence levels $h(X)$. The key property of $\rho$ which allows it to be used as a visualization aid is the following.
Proposition 4.8.
Suppose that $\phi$ and $\varphi$ satisfy the conditions of Lemma 4.6. Then the function $\rho$ as defined in (4.7) is nonnegative.
The proof of Proposition 4.8 (which can be found in Appendix 4.2) follows from applying Lemmas 4.1 and Lemma 4.6. The function $\rho$ shows how we trade off calibration with generalization at any confidence level $p$, and we can compute $\rho$ efficiently as it only requires estimating 1-D conditional expectations.
Our estimation of $\rho$ relies on Nadaraya-Watson kernel regression (Nadaraya, 1964; Watson, 1964) with a kernel $K$ and bandwidth hyperparameter $\sigma$ following the estimation approaches of Błasiok & Nakkiran (2023) and Chidambaram et al. (2024). We recall that given $n$ data points $(x_{i},y_{i})\in\mathbb{R}\times\mathbb{R}$ the kernel regression estimator of $\mathbb{E}[Y\mid X=x]$ is defined as:
Previous work (Zhang et al., 2020; Popordanoska et al., 2022; Chidambaram et al., 2024) has shown that we can obtain consistent estimators of $L^{p}$ ECE using (4.8). We extend these results to calibration errors derived from Bregman divergences. Due to the technical nature of some of the assumptions necessary on $K_{\sigma}$ for this result, we defer a proper presentation to Section 4.2 and present only an informal version of the result below. The proof relies on uniform convergence results for kernel regression applied to the individual terms in (4.1).
Theorem 4.9.
Constructing and reading calibration-sharpness diagrams. Using the above, we construct a diagram with three components: the calibration curve, the pointwise sharpness gap $\rho(h(X))$, and the density plot $\hat{p}_{h(X)}$. Example diagrams are shown in Figures 1 and 2 and interpreted in detail in Section 5. The calibration curve corresponds to $\mathbb{E}[\mathbbm{1}_{Y=c(x)}\mid h(x)]$ plotted at uniformly spaced values of $h(x)\in[0,1]$, and shows the conditional model accuracy vs. its confidence. A perfectly confidence-calibrated model satisfying $\mathbb{E}[\mathbbm{1}_{Y=c(x)}\mid h(x)]=h(x)$ is visualized as a dashed line, and any deviations from this dashed line can be interpreted as weaknesses in model calibration.
The sharpness gap is plotted as a band around the calibration curve, and the size of the band corresponds to (4.7) evaluated at each $h(x)$. A larger band is worse, as it means a bigger sharpness gap at a predicted confidence $h(x)$, which implies a worse generalization error at that particular confidence level. We also plot a kernel density estimate $\hat{p}_{h(X)}$ of the predicted confidences – this is intended to show where the model confidences concentrate. A density peak at $h(x)\approx 0.8$, for example, indicates that the model typically predicts around 80% confidence for its predicted class.
We point out that another common pitfall in several prior works is to omit this kind of density plot when discussing calibration, since it does not matter if a model is perfectly calibrated in a region of its confidence space that it rarely predicts. Lastly, we also overlay the estimated values of the confidence calibration error ($d_{\varphi,\mathrm{CAL}}$) and the full divergence ($d_{\phi,\mathrm{TOT}}$) on all plots, so as to give a sense of the full generalization error of the model as well as the part attributable to calibration. More precise estimation details for each component of the diagram are provided in Appendix D.
5 Experiments
To illustrate the benefits of calibration-sharpness diagrams, we first use them to revisit the experiments of Section 3 and then provide a direct comparison to the standard approach of reliability diagrams. For all of the experiments in this section, we focus on the ViT model of Section 3 for consistency and use Brier score/MSE since we will be comparing to ECE-based visualization methods in Section 5.2. We provide visualizations for other models and datasets in Appendix F. The kernel regressions estimates for the visualized components are computed using a Gaussian kernel with bandwidth $\sigma=0.05$; we discuss different choices of kernel and bandwidth in Appendix H. All of our experiments were done in PyTorch (Paszke et al., 2019) on a single A5000 GPU.
5.1 Revisiting MRR Experiments
(a) ViT (Baseline)
(b) ViT + TS
(c) ViT + HB
(d) ViT + IR
(e) ViT + MRR
Figure 1: Proposed calibration-sharpness diagrams using MSE/Brier score for the experiments of Section 3. Y-axes are not aligned here due to the large sharpness gap of MRR.
Figure 1 shows the calibration-sharpness diagrams (using $d_{\phi}(x,y)=\norm{x-y}^{2}$) for the ViT recalibration approaches in Table 1. The first thing to notice is that the calibration-sharpness diagram immediately reveals the weaknesses of both MRR and histogram binning - these methods using a small fixed set of predicted confidences (a single confidence in the case of MRR), and as a result incur large sharpness gaps as shown by the red shaded regions.
Additionally, when comparing the better recalibration approaches of temperature scaling and isotonic regression to the baseline model, we see that these algorithms have actually (perhaps counterintuitively) skewed the predicted confidence distributions towards even higher probabilities (as shown by the overlayed density plots of $\hat{p}_{h(X)}(h(x))$). Moreover, we see from the pointwise sharpness gap visualizations that the calibration improvements from isotonic regression (and to a lesser extent, temperature scaling) have come at the cost of some increased sharpness gaps. This can be expected from the fact that both methods led to larger predicted confidences, so mistakes incur larger errors.
5.2 Comparison to Reliability Diagrams
(a) HB, Calibration-Sharpness
(b) IR, Calibration-Sharpness
(c) HB, Reliability
(d) IR, Reliability
Figure 2: Comparing calibration-sharpness diagrams to reliability diagrams using histogram binning and isotonic regression. Y-axes of HB and IR are aligned for more direct comparison.
The standard approach taken by many empirical papers analyzing confidence calibration is to report reliability diagrams. These diagrams correspond to histogram binning estimators of the calibration error – namely, we sort the predicted confidences $h(x)$ into uniformly spaced buckets in the interval $[0,1]$ and plot a bar corresponding to the average accuracy in each bucket. Similar to the dashed line in our calibration-sharpness diagrams, perfectly confidence calibrated models would have bars that line up with the average confidence of each bucket (i.e. the height of the bar for the $[0,0.1]$ bucket should be somewhere in $[0,0.1]$).
Figure 2 shows reliability diagrams (using 15 bins, as was done in Guo et al. (2017)) for the histogram binning (HB) and isotonic regression (IR) recalibration approaches juxtaposed with axis-aligned calibration-sharpness diagrams from Figure 1. The reliability diagrams suggest that IR leads to better calibration than HB in the confidence range around $0.2$ and in $[0.5,0.9]$. However, these diagrams cannot be used to compare generalization and also do not give a sense of where predicted confidences concentrate.
On the other hand, the calibration-sharpness diagrams allow us to compare generalization while providing a more granular view of calibration performance as well. They show that (1) while the calibration of HB is worse than IR, the confidence region around $0.2$ is not a major problem for HB since there is no confidence density there and (2) that IR has better generalization performance (with respect to Brier score) due to smaller sharpness gaps.
Our work is not the first to point out these drawbacks of reliability diagrams for comparing calibration; the work of Błasiok & Nakkiran (2023) on SmoothECE also discusses such issues and puts forth a visualization that includes density plots combined with a kernel regression estimate of the calibration error as opposed to a binning estimate. However, their approach does not include a sharpness component, which in this instance is crucial for comparing HB and IR since this is where HB really loses out. We compare to the SmoothECE visualization approach in Appendix G.
6 Conclusion
The core tenets we hope to emphasize through both our theory and experiments are: report test accuracy and calibration error alongside the appropriate generalization metric, use a visualization of calibration that shows the density of predicted confidences, and be aware of whether improvements in calibration in some regions of predicted confidence come at the cost of sharpness/generalization. We believe there is still much work to be done on addressing how to compare and visualize calibration/generalization metrics, and we hope future work continues to push in this direction.
Ethics Statement
While calibration has become an increasingly important topic given the use of machine learning models for decision-making, our work concerns only the theory and practice of reporting calibration. As a result, we do not anticipate any negative societal impacts or significant ethical considerations as consequences of this work.
Reproducibility Statement
Full proofs of all results in the main paper can be found in the appendices following the references. The code necessary to recreate the experiments in the main paper as well as the experiments in the appendices can be found in the supplementary material, as well as in the Python package sharpcal.
Acknowledgments
Rong Ge and Muthu Chidambaram were supported by NSF Award DMS-2031849 and CCF-1845171 (CAREER) during the completion of this work.
Appendix A Background Results
See 4.1